OpenStudy (anonymous):

x^(2/3)-2x^(1/3)-3=0 x=___ simplify, use radicals as needed, express complex numbers in terms of i

OpenStudy (anonymous):

x = 27, do you need the work?

OpenStudy (anonymous):

no... thnk u!!!

OpenStudy (anonymous):

x^1/3 = a x^2/3 = [x^1/3]^2 = x^2 x^(2/3)-2x^(1/3)-3=0 becomes a^2 -2a -3 =0 We will solve first for a a^2 -3a +a -3 =0 a(a-3) +1 (a-3) = 0 (a-3) (a+1) =0 a-3 =0 means a = 3 a+1 =0 means a = -1 Plug in x^1/3 = a x =a^3 a =3 means x = 27 a=-1 means x = -1